Use the information provided below to answer each of the following questions independently:

2.1

Calculate the break-even quantity.

(2 Marks)

2.2

Calculate the sales value required to achieve a net profit of R150 000, using the marginal income ratio.

(4 Marks)

2.3

Determine the selling price per unit if a net profit of R624 600 is desired.

(4 Marks)

INFORMATION

The following information was extracted from the budget of Mary’s Manufacturers for the period July to December 2021:

1.

Total production and sales

2 300 units

2.

Selling price per unit

R200

3.

Variable manufacturing costs per unit:

Direct materials

R60

Direct labour

R40

Overheads

R20

4.

Fixed manufacturing overheads

R200 000

5.

Other costs:

Fixed marketing and administrative costs

R100 000

Sales commission

10% of sales


1
Expert's answer
2021-06-07T11:28:25-0400

2.1

Margin profit:

"200*0.9-60-40-20=60"

the break-even quantity:

"\frac{200\ 000+100\ 000}{60}=5\ 000" units

2.2

the  sales value required to achieve a net profit of R150 000:

"\frac{200\ 000+100\ 000+150\ 000}{60}=7\ 500" units

2.3

"\frac{200\ 000+100\ 000+624\ 600}{x*0.9-60-40-20}=2\ 300"

"x=580" R